General de la Roverie

Greg Hudson

October 14, 1994

General de la Roverie had some important existential features which made Germaldi at the end of the movie, though imprisoned and executed, seem more powerful and free than the Germaldi at the beginning.

Most obvious is Germaldi’s willingness to take responsibility for his actions. At the beginning of the movie, he is constantly in a bind where he needs more money and referring to “hard times” as an excuse for his situation. Against his own self-interest, he gambles away the money he cons from people. When Germaldi is caught, and asked by the captain whether he wants to be tried by a military court or go back to his regular life, he says, “How can I make that decision? If I go before a military court, I’ll be shot,” completely discarding the possibility of accepting punishment for his actions.

The military prison illustrates the existential claim that the world is absurd. In an environment where the actions of the prisoners are highly restricted and their fates are decided by the politics of the German military, they nevertheless labor futilely to organize themselves, even committing suicide to avoid releasing information. But these prisoners manage to transcend the absurdity of their situation, despite the futility of their labors, by avoiding hope. Banchelli mentions to Germaldi that he has already been sentenced to death, but that they just haven’t gotten around to it yet; he has no hope, but still acts as a go-between between Germaldi and Fermazio.

Seeing this, Germaldi learns to transcend his own situation, and at the end of the movie, takes responsibility for his actions according to what he cares about. Rather than allow the Germans to discover Fermazio’s identity (by bribery or by torture), he allows himself to be executed by a firing squad and become a martyr of the resistance. In so doing, he illustrates another existentialist claim: that even though he is not, in fact, General de la Roverie, his actions can make him effectively that general. Like the General, he acts as an inspiration to the other prisoners, and like the general, he has knowledge whose secrecy is crucial to the resistance.